OpenStudy (anonymous):

help simplify and show work please! (4+i)^2

OpenStudy (anonymous):

i for imaginary number?

OpenStudy (anonymous):

yes!

OpenStudy (cwrw238):

16 + 8i - 1 = 15 + 8i

OpenStudy (cwrw238):

i ^ 2 = -1 ( by definition)

OpenStudy (anonymous):

(4+i)(4+i) = 16+8i+i^2 Since i^2=-1 you get: 15+8i Hope this helps!?! :) //Ali

OpenStudy (anonymous):

thanks everyone!